The total cost to rent costumes for the school play is $180.00. Since the boys' costumes are more elaorate, they cost $9.00 each, while the girls' costumes cost $4.50 each.
a. Write the equation that represents the scenario. Let x represent the cost of a boys' costume and y represent the cost of a girls' costume.
b. Solve the equation for y.

Answer by checkley77(12569)

9X+4.5Y=180
4.5Y=180-9X
Y=(180-9X)/4.5
Y=40-2X
Y=2(20-X) ANS.